Julie went to the post office and bought both $0.49 stamps and $0.34 postcards for her office's bills. She spent $75.80. The number of stamps was 20 more than twice the number of postcards. How many of each did she buy
Answer by stanbon(75887)

Julie went to the post office and bought both $0.49 stamps and $0.34 postcards for her office's bills. She spent $75.80. The number of stamps was 20 more than twice the number of postcards. How many of each did she buy
--------
Equations:
s = 2p + 20
0.49s + 0.34p = 75.80
---------------------------------
49s - 98p = 49*20
49s + 34p = 7580
----------------------------
Subtract and solve for "p"::
132p = 6600
----
p = 50 (# of postcards)
s = 2p+20 = 120 (# of stamps)
------------
Cheers,
Stan H.
-----------